4(x-1)+6 = 3x+1-2x
Answer by KneesAndToes(13)

4(x-1)+6 = 3x+1-2x ... (In this problem, it's easiest to multiply out first)


4x-4+6 = 3x+1-2x ... (Add like terms on each side)


4x+2 = 1x+1 ... (Move unknowns to one side, knowns to the other)


4x-1x = 1-2 ... (Simplify)


3x = -1 ... (Divide each side by 3)


x = -1/3


Check your answer:
4((-1/3)-1)+6 = 3(-1/3)+1-2(-1/3)
2/3 = 2/3